Please, solve all questions with clear steps, or please let someone else do it.

4. Wild E. Coyote fires an Acme rocket with an initial velocity of 35.0 m/s at an angle of 38 above the horizontal ground. Find the following (a) The rockets time in the air (b) The rockets maximum height (c) The distance the rocket travels horizontally
Answer #1

Initial velocity, u= 35.0m/s at an angle \theta= 38^\circ above horizontal

then Its horizontal component, u_x= ucos\theta= (35) cos38^\circ= 27.58 m/s

its vertical component, u_y= usin\theta= (35) sin38^\circ= 21.548 m/s

a). We assume air firction to be negligible, then motion will be completly uniform and thus

its time of ascent= time of descent=t

and Total time of flight or Total time in air will be, T= t+t

T=2t............(1)

Now considering its vertical motion, at highest point its final velocity will be zero.

Hence, vf=0m/s

its vertical component of initial velocity, u_y= 21.548 m/s

acceleration, ay= -g=-9.8m/s2 ( As gravity is against its vertical motion)

time= t=T/2

Now using equation of motion,

v_f= u_y+a_yt

0= (21.548)+(-9.8)(\frac{T}{2}) (using equation 1)

T= \frac{21.548 \times 2}{9.8}= 4.397 s(ANS)

b). Again considering its vertical motion and assuming that it maximum height is H, then using equations of motion,

  v_f^2-u_y^2= 2a_yH

(0)^2-(21.548)^2= 2(-9.8)H

  H= 23.689 m(ANS)

c). Also In time "T", max horizontal distance projectile will cover without air friction is,

Horizontal range= ( Horizontal component of velocity)(Time of flight) (Since no acceleration acts horizontally)

R= u_xT= 27.58 \times 4.397= 121.269 m(ANS)